Random sample of size N = 2 are drawn from a finite population consisting of


the numbers 1, 3, 8, 9.


Find the following:


Mean, Variance and Standard deviation of the Population.


Mean, Variance and Standard deviation of the Sample means.

1
Expert's answer
2022-04-16T04:12:56-0400

The population mean:

"\mu=\cfrac{1+3+8+9}{4}=5.25."


The population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\begin{Bmatrix}
 1-5.25,3-5.25,8-5.25,9-5.25
\end{Bmatrix}="

"=\begin{Bmatrix}
-4.25, -2.25, 2.75, 3.75
\end{Bmatrix},"

"\sigma^2=(-4.25)^2\cdot \cfrac{1}{4}+(-2.25)^2\cdot \cfrac{1}{4}+2.75^2\cdot \cfrac{1}{4}+\\
+3.75^2\cdot \cfrac{1}{4}=11.189."


The population standard deviation:

"\sigma=\sqrt{11.189}=3.345."


The mean of the sampling distribution of sample means:

"\mu_{\bar x} =\mu=5.25."


The variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{11.189}{2}=5.594."


The standard deviation of the sampling distribution of sample means:

"\sigma_{\bar x}=\cfrac{\sigma}{\sqrt n}=\cfrac{3.345}{\sqrt 2}=2.365."